OpenStudy (roxy.girl):

Solve for B. 13+6-3x=4b

OpenStudy (roxy.girl):

@Qwertty123

Mehek (mehek14):

Combine like terms 13 + 6 = ?

OpenStudy (qwertty123):

Yes you would want to Combine like terms: 13 + 6 =?

OpenStudy (roxy.girl):

19

Mehek (mehek14):

alright so we have 19 - 3x = 4b

OpenStudy (qwertty123):

Move all terms containing b to the left, all other terms to the right. Now what do you do to get b by itself?

OpenStudy (roxy.girl):

add -4?

OpenStudy (qwertty123):

yes, you now add '-4b' to each side of the equation. 19 + -4b + -3x = 4b + -4b

OpenStudy (qwertty123):

Combine like terms: 4b + -4b =?

OpenStudy (roxy.girl):

0

OpenStudy (qwertty123):

yes, 19 + -4b + -3x = 0

OpenStudy (roxy.girl):

add -19?

OpenStudy (qwertty123):

Yes! Add '-19' to each side of the equation. 19 + -4b + -19 + -3x = 0 + -19 Reorder the terms: 19 + -19 + -4b + -3x = 0 + -19 Combine like terms: 19 + -19 =?

OpenStudy (roxy.girl):

0

OpenStudy (qwertty123):

yes okay so it looks like this after we... Combine like terms: 19 + -19 = 0 0 + -4b + -3x = 0 + -19 -4b + -3x = 0 + -19 Combine like terms: 0 + -19 = ?

OpenStudy (roxy.girl):

-19

OpenStudy (roxy.girl):

so now would you add 3?

OpenStudy (qwertty123):

yes! Now Add '3x' to each side of the equation. -4b + -3x + 3x = -19 + 3x Combine like terms: -3x + 3x =?

OpenStudy (roxy.girl):

0

OpenStudy (qwertty123):

yes -4b + 0 = -19 + 3x -4b = -19 + 3x

OpenStudy (roxy.girl):

DIVIDE -4?

OpenStudy (qwertty123):

yes, Divide each side by '-4'. What do you get?

OpenStudy (mathmale):

Once you have a solution, you might want to determine whether there's a quicker way to solve this equation for b. If 13+6-3x=4b, then 19-3x=4b. Must enlose that "19-3x) within parentheses. Then divide that quantity (within parentheses) by 4. Done.

OpenStudy (qwertty123):

Thanks Mathmale! :3

OpenStudy (roxy.girl):

4.75 + -0.75x

OpenStudy (qwertty123):

yes